Find two consecutive odd integers such that the difference of their squares is 160.
Answer by mananth(16946)

let the two consecutive odd integers be (x+2) & x
(x+2)^2-x^2=160

x^2+4x+4-x^2=160
4x+4=160
4x=156
x=39
39 & 41